# TabKAN: Advancing Tabular Data Analysis using Kolmograv-Arnold Network

### v1

@@ Abstract
Tabular data analysis presents unique challenges due to its heterogeneous feature types, missing values, and complex interactions. While traditional machine learning methods, such as gradient boosting, often outperform deep learning approaches, recent advancements in neural architectures offer promising alternatives. This paper introduces TabKAN, a novel framework that advances tabular data modeling using Kolmogorov-Arnold Networks (KANs). Unlike conventional deep learning models, KANs leverage learnable activation functions on edges, enhancing both interpretability and training efficiency. Our contributions include: (1) the introduction of modular KAN-based architectures tailored for tabular data analysis, (2) the development of a transfer learning framework for KAN models, enabling effective knowledge transfer between domains, (3) the development of model-specific interpretability for tabular data learning, reducing reliance on post hoc and model-agnostic analysis, and (4) comprehensive evaluation of vanilla supervised learning across binary and multi-class classification tasks. Through extensive benchmarking on diverse public datasets, TabKAN demonstrates superior performance in supervised learning while significantly outperforming classical and Transformer-based models in transfer learning scenarios. Our findings highlight the advantage of KAN-based architectures in efficiently transferring knowledge across domains, bridging the gap between traditional machine learning and deep learning for structured data.
@@ 1 Introduction
Tabular data is a fundamental form of structured data prevalent across diverse domains, ranging from healthcare and finance to e-commerce and scientific research. Analyzing such data efficiently and accurately is critical for deriving actionable insights and advancing decision-making processes. Despite the significant progress achieved in deep learning, particularly in image and text processing, tabular data analysis remains a challenging domain due to its unique characteristics, such as heterogeneous feature types, missing values, and complex feature interactions. While traditional machine learning methods, such as tree-based models, have frequently demonstrated superior performance compared to deep learning approaches in this domain, the development of novel architectures specifically designed for tabular data remains crucial.
In recent years, the potential of neural networks for tabular data modeling has gained significant attention. Among these, Kolmogorov-Arnold Networks have emerged as a promising alternative to traditional Multi-Layer Perceptrons, addressing key challenges such as interpretability, training efficiency and robustness. Unlike MLPs, which rely on fixed activation functions at nodes, KANs feature learnable activation functions on edges, enabling more flexible and efficient modeling of complex feature interactions. Inspired by the Kolmogorov-Arnold representation theorem, which expresses any multivariate continuous function as a composition of univariate functions and a summation operator, KANs provide a structured framework for tabular data analysis.
This paper introduces TabKAN, an encompassing framework for numerical and categorical feature encoding through a KAN-based Module specifically designed for tabular data modeling using various KAN-based architectures, including spl-KAN [1], ChebyshevKAN [2], RKAN [3], Fourier KAN [4], fKAN [5] and Fast-KAN [6]. Due to the inherent diversity and complexity of patterns present in tabular datasets, TabKAN leverages multiple KAN variants, enabling flexible adaptation and enhanced representation capabilities tailored to different data characteristics. The primary novelties of this study are:
  • The introduction of modular KAN-based architectures tailored for tabular data analysis
  • The development of a transfer learning framework for KAN models, enabling effective knowledge transfer between domains
  • The creation of model-specific interpretability methods for tabular data learning, reducing reliance on post hoc and model-agnostic analysis
@@
TabKAN demonstrates stable and significantly improved performance in supervised learning and transfer learning tasks outperforming baseline models in binary and multi-class classification across diverse public datasets. By applying KAN principles, TabKAN unifies traditional machine learning and deep learning, providing a robust, interpretable, and efficient solution for tabular data modeling.
@@ 2 Related Work
Existing methods for tabular learning face multiple obstacles, such as mismatched feature sets between training and testing, limited or missing labels, and the potential emergence of new features over time [7]. It can be categorized as:
Classic Machine Learning Models. Early techniques rely on parametric or non-parametric strategies like K-Nearest Neighbors (KNN), Gradient Boosting, Decision Trees, and Logistic Regression [8]. Popular models include Logistic Regression (LR), XGBoost [9, 10], and MLP. A notable extension is the self-normalizing neural network (SNN) [11], which employs scaled exponential linear units (SELU) to maintain neuron activations at zero mean and unit variance. While SNNs are straightforward and effective, they can struggle with complex, high-dimensional data, prompting the development of more advanced neural architectures.
Deep Learning-Based Supervised Models. Building on Transformer architectures, methods such as AutoInt [12] use self-attention to learn feature importance, while TransTab [13] extends Transformers to handle partially overlapping columns across multiple tables, enabling tasks like transfer learning, incremental feature learning, and zero-shot inference. Similarly, TabTransformer [14] applies self-attention to enhance feature embeddings, achieving robust performance even with missing data. SAINT [15] introduces a hybrid attention scheme at both row and column levels, pairing it with inter-sample attention and contrastive pre-training, and surpasses gradient boosting models including XGBoost [9], CatBoost [16], and LightGBM [17] on various benchmarks. Other innovations include TabRet [18], which implements a retokenization step for previously unseen columns, and XTab [19], which enables cross-table pretraining in a federated learning setup, managing heterogeneous column types and numbers. TabCBM [20] provides concept-based explanations that allow human oversight, thereby striking a balance between strong predictive accuracy and transparent, understandable reasoning. TabPFN [21] is a pretrained Transformer-based model that performs zero-shot classification on tabular data by leveraging meta-learning, eliminating the need for task-specific training. TabMap [22] transforms tabular data into 2D topographic maps that spatially encode feature relationships and preserve values as pixel intensities, allowing convolutional neural networks to efficiently extract association patterns and outperform other deep learning-based supervised models. TabSAL [23] employs lightweight language models to generate privacy-free synthetic tabular data when raw data cannot be shared due to privacy risks. TabMixer [24] is an enhanced MLP-mixer-based architecture that captures both sample-wise and feature-wise interactions using a self-attention mechanism.
@@ 3.1 Data Preprocessing
In order to handle missing values and class imbalance we used the preprocessing method in [24]. Let the input variable space as $\mathcal{D}ata\in\{\mathbb{R}\cup\mathbb{C}\cup\mathbb{B}\cup\varnothing\}$, where $\mathbb{R}$, $\mathbb{C}$, and $\mathbb{B}$ represent the domains of numerical, categorical, and binary data, respectively. After preprocessing block, we have $\{\mathcal{X},\mathcal{Y}\}$ denote a feature-target pair, where $x$ represents numerical features, and $y$ is integer number for classification tasks. $\{\mathcal{Y}\}$ can have dimension of one for binary classification or $M$ for multi-class classification.
In most tables, the data consists of both continuous numerical and categorical variables. We preprocess the categorical features by transforming them into one-hot vectors. After the preprocessing step, the data is represented as $n\times m$ matrix, where all entries are numerical values.
3.2 Kolmogorov–Arnold Networks (KANs)
@@
$\text{KAN}(\mathbf{x}_{0})=\bigl{(}\Phi_{L-1}\circ\cdots\circ\Phi_{0}\bigr{)}\,\mathbf{x}_{0}$
@@
$f(\mathbf{x})=\text{KAN}(\mathbf{x})$
@@
This construction mirrors that of Multi-Layer Perceptrons (MLPs), where a single layer, an affine transformation followed by a nonlinearity, is stacked to increase model capacity. Similarly, we define a KAN layer and compose multiple such layers to build a deeper, more expressive model.
@@
$\tilde{x}_{\ell,j,i}=\phi_{\ell,j,i}\bigl{(}x_{\ell,i}\bigr{)}$
@@
$\mathbf{x}_{\ell+1}=\begin{pmatrix}\phi_{\ell,1,1}(\cdot)&\cdots&\phi_{\ell,1,n_{\ell}}(\cdot)\\ \vdots&\ddots&\vdots\\ \phi_{\ell,n_{\ell+1},1}(\cdot)&\cdots&\phi_{\ell,n_{\ell+1},n_{\ell}}(\cdot)\end{pmatrix}\mathbf{x}_{\ell}$
@@
3.3 General KAN Layer
To define a KAN layer with $n_{\text{in}}$-dimensional inputs and $n_{\text{out}}$-dimensional outputs, consider a matrix of one-dimensional functions:
@@ 3.3 General KAN Layer
where each $\phi_{q,p}$ has trainable parameters. In the original Kolmogorov–Arnold theorem, the “inner” functions form a KAN layer with $n_{\text{in}}=n$ and $n_{\text{out}}=2n+1$, while the “outer” functions form another KAN layer with $n_{\text{in}}=2n+1$ and $n_{\text{out}}=1$. Since these operations are all differentiable, we train KANs through standard backpropagation.
3.4 Chebyshev Kolmogorov-Arnold Network Layer (ChebyKAN)
The ChebyKANLayer [2] employs Chebyshev polynomials of the first kind, $\{T_{k}(x)\}_{k=0}^{d}$, to approximate nonlinear functions with fewer parameters than traditional MLPs. First, the input $\mathbf{x}\in\mathbb{R}^{n}$ is normalized to $[-1,1]$ via the hyperbolic tangent function:
$\tilde{\mathbf{x}}=\tanh(\mathbf{x})$
@@ 3.4 Chebyshev Kolmogorov-Arnold Network Layer (ChebyKAN)
Let $\Theta\in\mathbb{R}^{n\times m\times(d+1)}$ be the trainable coefficient tensor for $n$ input features, $m$ outputs, and polynomial degree $d+1$. The layer output is computed via Einstein summation:
$y_{bo}=\sum_{i=1}^{n}\sum_{k=0}^{d}T_{bik}\,\Theta_{iok}$ (1)
@@
3.5 Fast Kolmogorov-Arnold Network (Fast KAN)
@@
$f(x)\;=\;\sum_{i=1}^{N}w_{i}\,\phi\bigl{(}\|\mathbf{x}-\mathbf{c}_{i}\|\bigr{)}$
@@
$\phi(r)\;=\;\exp\left(-\tfrac{r^{2}}{2h^{2}}\right)$
@@
$\mathbf{y}\;=\;f_{L}\circ f_{L-1}\circ\cdots\circ f_{1}(\mathbf{x})$
We optimize the number of layers $L$ and the width vector $\mathbf{w}=(w_{1},\ldots,w_{L})$ as hyperparameters to adapt the Fast KAN architecture to different datasets. This includes both the number of layers and the neurons per layer. Additionally, the parameters of the RBF activation functions (e.g., centers and bandwidths) can be tuned or learned. To efficiently explore architectural configurations, we employ the automated hyperparameter optimization framework Optuna [27], selecting the settings that maximize validation performance (F1 Score).
3.6 Padé Rational Kolmogorov-Arnold Network (PadéRKAN)
@@
$R(x)=\frac{P_{q}(x)}{Q_{k}(x)}=\frac{\sum_{i=0}^{q}a_{i}\,x^{i}}{\sum_{j=0}^{k}b_{j}\,x^{j}}$
In each PadéRKAN layer, this rational form serves as the activation function, enabling the model to capture asymptotic behavior and abrupt transitions more accurately. Specifically, for an input $\mathbf{x}\in\mathbb{R}^{d}$, the layer outputs
$\mathbf{y}=\frac{\sum_{i=0}^{q}\theta_{i}\,P_{i}(\mathbf{x})}{\sum_{j=0}^{k}\theta_{j}\,Q_{j}(\mathbf{x})}$
@@
By tuning the network depth, layer widths, and polynomial degrees $(q,k)$ with Optuna, PadéRKAN adapts to various datasets and task complexities, while the L-BFGS optimizer minimizes the cross-entropy loss. Due to its rational formulation, PadéRKAN is capable of approximating functions characterized by sharp transitions and asymptotic behaviors, thereby providing enhanced modeling flexibility relative to purely polynomial-based KAN variants.
3.7 Fourier Kolmogorov-Arnold Network (Fourier KAN)
Fourier KAN [28] leverages a Fourier series expansion to capture both low- and high-frequency components in tabular or structured data. Given an input vector $\mathbf{x}\in\mathbb{R}^{d}$, the transformation function $\phi_{F}(\mathbf{x})$ introduces sine and cosine terms up to a grid size $g$, thus enabling the network to approximate highly complex or oscillatory functions. Formally,
$\phi_{F}(\mathbf{x})=\sum_{i=1}^{d}\sum_{k=1}^{g}\bigl{(}a_{ik}\cos(k\,x_{i})+b_{ik}\sin(k\,x_{i})\bigr{)}$
@@ 3.7 Fourier Kolmogorov-Arnold Network (Fourier KAN)
$y=\sum_{i=1}^{d}\sum_{k=1}^{g}\Bigl{(}W_{ik}^{(c)}\cos(k\,x_{i})+W_{ik}^{(s)}\sin(k\,x_{i})\Bigr{)}+b$ (2)
where $W_{ik}^{(c)}$ and $W_{ik}^{(s)}$ are learnable weights for the cosine and sine terms, respectively, and $b$ is a bias. By capitalizing on the orthogonality of trigonometric functions, Fourier KAN often converges faster than traditional MLPs and polynomial-based KANs while reducing overfitting. In practice, tuning the grid size $g$ is critical to accommodate the desired frequency range and optimize performance on diverse datasets.
3.8 Fractional Kolmogorov-Arnold Network (fKAN)
The Fractional Kolmogorov-Arnold Network (fKAN) [29] incorporates fractional-order Jacobi polynomials into the Kolmogorov-Arnold framework to enhance expressiveness and adaptability. Each layer of fKAN uses a Fractional Jacobi Neural Block (fJNB), which introduces a trainable fractional parameter $\nu$ to adjust the polynomial basis dynamically. For an input $\mathbf{x}\in\mathbb{R}^{d}$, the fractional Jacobi polynomial $J_{n}^{(\alpha,\beta,\nu)}(x)$ is given by:
$J_{n}^{(\alpha,\beta)}(x^{\nu})=\frac{(\alpha+1)_{n}}{n!}\sum_{k=0}^{n}\binom{n}{k}\frac{(\beta+1)_{n-k}}{(\alpha+\beta+1)_{n-k}}\left(\frac{x^{\nu}-1}{2}\right)^{k}\left(\frac{x^{\nu}+1}{2}\right)^{n-k}$
where $(\alpha,\beta)>-1$ determine the shape of the polynomial, and $B(\cdot,\cdot)$ denotes the Beta function.
Within fKAN, each layer applies a linear transformation followed by a fractional Jacobi activation, enabling the model to capture subtle data patterns. Hyperparameters such as network depth, layer widths, and the fractional order $\nu$ are tuned with Optuna, and training is performed using the L-BFGS optimizer to minimize cross-entropy loss. By allowing the polynomial order to be learned, fKAN can adapt its functional complexity to a wide range of datasets, often yielding improved generalization as evidenced by metrics like F1 score and ROC-AUC.
3.9 Rational Kolmogorov-Arnold Network (RKAN)

### v2

# TabKAN: Advancing Tabular Data Analysis using Kolmogorov-Arnold Network

@@ Abstract
Tabular data analysis presents unique challenges due to its heterogeneous feature types, missing values, and complex interactions. While traditional machine learning methods, such as gradient boosting, often outperform deep learning approaches, recent advancements in neural architectures offer promising alternatives. This paper introduces TabKAN, a novel framework that advances tabular data modeling using Kolmogorov-Arnold Networks (KANs). Unlike conventional deep learning models, KANs leverge learnable activation functions on edges, which improve both interpretability and training efficiency. Our contributions include: (1) the introduction of modular KAN-based architectures for tabular data analysis, (2) the development of a transfer learning framework for KAN models that supports knowledge transfer between domains, (3) the development of model-specific interpretability for tabular data learning, which reduces dependence on post hoc and model-agnostic analysis, and (4) comprehensive evaluation of vanilla supervised learning across binary and multi-class classification tasks. Through extensive benchmarking on diverse public datasets, TabKAN demonstrates superior performance in supervised learning while significantly outperforming classical and Transformer-based models in transfer learning scenarios. Our findings highlight the advantage of KAN-based architectures in transferring knowledge across domains and narrowing the gap between traditional machine learning and deep learning for structured data.
@@ 1 Introduction
Tabular data is a fundamental form of structured data commonly observed across diverse domains, including healthcare, finance, e-commerce, and scientific research. Analyzing such data efficiently and accurately is critical for deriving actionable insights and improving decision-making processes. Despite significant progress in deep learning, particularly in image and text processing, tabular data analysis remains a challenging area due to its unique characteristics, such as heterogeneous feature types, missing values, and complex feature interactions. While traditional machine learning methods, such as tree-based models, have often shown stronger performance than deep learning approaches in this domain, the development of novel architectures specific to tabular data remains important.
In recent years, the potential of neural networks for tabular data modeling has gained increasing attention. Among these, Kolmogorov-Arnold Networks (KANs) have emerged as a promising alternative to traditional Multi-Layer Perceptrons (MLPs), addressing key challenges such as interpretability, training efficiency, and robustness. Unlike MLPs, which rely on fixed activation functions at nodes, KANs use learnable activation functions on edges that permits more flexible and efficient modeling of complex feature interactions. Inspired by the Kolmogorov-Arnold representation theorem, which expresses any multivariate continuous function as a composition of univariate functions and a summation operator, KANs offer a structured framework for tabular data analysis.
This paper introduces TabKAN, a framework for numerical and categorical feature encoding through a KAN-based module developed specifically for tabular data modeling using various KAN-based architectures, including spline-KAN [1], ChebyshevKAN [2], RKAN [3], Fourier KAN [4], fKAN [5], and Fast-KAN [6]. Due to the diversity and complexity of patterns in tabular datasets, TabKAN uses multiple KAN variants to allow flexible adaptation and improved representation capacity suited to different data characteristics. The primary novelties of this study are:
  • The introduction of modular KAN-based architectures for tabular data analysis
  • The development of a transfer learning framework for KAN models that supports knowledge transfer between domains
  • The creation of model-specific interpretability methods for tabular data learning that reduce reliance on post hoc analysis
@@
TabKAN demonstrates stable and significantly improved performance in supervised learning and transfer learning tasks, outperforming baseline models in binary and multi-class classification across diverse public datasets. By applying KAN principles, TabKAN unifies traditional machine learning and deep learning, providing a robust, interpretable, and efficient solution for tabular data modeling.
Unlike traditional regression-based approaches used to explore factors affecting health-related quality of life [7], TabKAN offers a novel, attention-based method suitable for modeling complex interactions in healthcare data.
@@ 2 Related Work
Existing methods for tabular learning face multiple obstacles, such as mismatched feature sets between training and testing, limited or missing labels, and the potential emergence of new features over time [8]. These methods can be categorized as:
Classic Machine Learning Models. Early techniques rely on parametric or non-parametric strategies like K-Nearest Neighbors (KNN), Gradient Boosting, Decision Trees, and Logistic Regression [9]. Popular models include Logistic Regression (LR), XGBoost [10, 11], and MLP. A notable extension is the self-normalizing neural network (SNN) [12], which uses scaled exponential linear units (SELU) to maintain neuron activations at zero mean and unit variance. While SNNs are simple and effective, they can fail on complex, high-dimensional data, which has led to the proposal of more advanced neural architectures.
Deep Learning-Based Supervised Models. Building on Transformer architectures, methods such as AutoInt [13] apply self-attention to learn feature importance, while TransTab [14] extends Transformers to handle partially overlapping columns across multiple tables. This extension supports tasks such as transfer learning, incremental feature learning, and zero-shot inference. TabTransformer [15] applies self-attention to improve feature embeddings and achieves strong performance even with missing data. SAINT [16] introduces hybrid attention at both row and column levels, pairs it with inter-sample attention and contrastive pre-training, and outperforms gradient boosting models including XGBoost [10], CatBoost [17], and LightGBM [18] on several benchmarks.
Other innovations include TabRet [19], which implements a retokenization step for previously unseen columns, and XTab [20], which enables cross-table pretraining in a federated learning setup, handling heterogeneous column types and numbers. TabCBM [21] introduces concept-based explanations that support human oversight, balancing predictive accuracy and interpretability. TabPFN [22] is a pretrained Transformer that performs zero-shot classification on tabular data through meta-learning, without requiring task-specific training. TabMap [23] transforms tabular data into 2D topographic maps that encode feature relationships spatially and preserve values as pixel intensities. This structure allows convolutional networks to detect association patterns efficiently and outperforms other deep learning-based supervised models. TabSAL [24] employs lightweight language models to generate privacy-free synthetic tabular data when raw data cannot be shared due to privacy concerns. TabMixer [25] builds on the MLP-mixer framework and captures both sample-wise and feature-wise interactions through a self-attention mechanism.
@@ 3.1 Data Preprocessing
To address missing values and class imbalance, we adopted the preprocessing strategy introduced in [25]. Let the input variable space be defined as $\mathcal{D}ata\in\{\mathbb{R}\cup\mathbb{C}\cup\mathbb{B}\cup\varnothing\}$, where $\mathbb{R}$, $\mathbb{C}$, and $\mathbb{B}$ denote the domains of numerical, categorical, and binary data, respectively. After the preprocessing block, we denote the resulting feature-target pair as $\{\mathcal{X},\mathcal{Y}\}$, where $x$ contains numerical features, and $y$ is an integer used in classification tasks. The label set $\{\mathcal{Y}\}$ may have dimension one for binary classification or $M$ for multi-class classification.
Most tabular datasets contain both continuous numerical and categorical variables. We preprocess the categorical features by converting them into one-hot vectors. After preprocessing, the data is organized as an $n\times m$ matrix with purely numerical entries.
3.2 Kolmogorov-Arnold Networks (KANs)
3.2.1 Spline Kolmogorov-Arnold Network
@@
$\text{KAN}(\mathbf{x}_{0})=\bigl{(}\Phi_{L-1}\circ\cdots\circ\Phi_{0}\bigr{)}\,\mathbf{x}_{0}.$
@@
$f(\mathbf{x})=\text{KAN}(\mathbf{x}),$
@@
This construction mirrors that of MLPs, where a single layer, an affine transformation followed by a nonlinearity, is stacked to increase model capacity. Similarly, we define a KAN layer and compose multiple such layers to build a deeper, more expressive model.
@@
$\tilde{x}_{\ell,j,i}=\phi_{\ell,j,i}\bigl{(}x_{\ell,i}\bigr{)}.$
@@
$\mathbf{x}_{\ell+1}=\begin{pmatrix}\phi_{\ell,1,1}(\cdot)&\cdots&\phi_{\ell,1,n_{\ell}}(\cdot)\\ \vdots&\ddots&\vdots\\ \phi_{\ell,n_{\ell+1},1}(\cdot)&\cdots&\phi_{\ell,n_{\ell+1},n_{\ell}}(\cdot)\end{pmatrix}\mathbf{x}_{\ell},$
@@
To define a spline KAN layer with $n_{\text{in}}$-dimensional inputs and $n_{\text{out}}$-dimensional outputs, consider a matrix of one-dimensional functions:
@@
where each $\phi_{q,p}$ has trainable parameters. In the original Kolmogorov-Arnold theorem, the “inner” functions form a KAN layer with $n_{\text{in}}=n$ and $n_{\text{out}}=2n+1$, while the “outer” functions form another KAN layer with $n_{\text{in}}=2n+1$ and $n_{\text{out}}=1$. Since these operations are all differentiable, we train KANs through standard backpropagation.
3.2.2 Chebyshev Kolmogorov-Arnold Network Layer (ChebyKAN)
The ChebyKAN [2] employs Chebyshev polynomials of the first kind, $\{T_{k}(x)\}_{k=0}^{d}$, to approximate nonlinear functions with fewer parameters than traditional MLPs. First, the input $\mathbf{x}\in\mathbb{R}^{n}$ is normalized to $[-1,1]$ via the hyperbolic tangent function:
$\tilde{\mathbf{x}}=\tanh(\mathbf{x}).$
@@ 3.2.2 Chebyshev Kolmogorov-Arnold Network Layer (ChebyKAN)
Let $\Theta\in\mathbb{R}^{n\times m\times(d+1)}$ be the trainable coefficient tensor for $n$ input features, $m$ outputs, and polynomial degree $d+1$. The output of the ChebyKAN layer is computed via Einstein summation:
$y_{bo}=\sum_{i=1}^{n}\sum_{k=0}^{d}T_{bik}\,\Theta_{iok},$ (1)
@@
3.2.3 Fast Kolmogorov-Arnold Network (Fast KAN)
@@
$f(x)\;=\;\sum_{i=1}^{N}w_{i}\,\phi\bigl{(}\|\mathbf{x}-\mathbf{c}_{i}\|\bigr{)},$
@@
$\phi(r)\;=\;\exp\left(-\tfrac{r^{2}}{2h^{2}}\right),$
@@
$\mathbf{y}\;=\;f_{L}\circ f_{L-1}\circ\cdots\circ f_{1}(\mathbf{x}).$
3.2.4 Padé Rational Kolmogorov-Arnold Network (PadéRKAN)
@@
$R(x)=\frac{P_{q}(x)}{Q_{k}(x)}=\frac{\sum_{i=0}^{q}a_{i}\,x^{i}}{\sum_{j=0}^{k}b_{j}\,x^{j}}.$
In each PadéRKAN layer, this rational form acts as the activation function. This structure allows the model to capture asymptotic behavior and abrupt transitions with greater precision. Specifically, for an input $\mathbf{x}\in\mathbb{R}^{d}$, the layer outputs
$\mathbf{y}=\frac{\sum_{i=0}^{q}\theta_{i}\,P_{i}(\mathbf{x})}{\sum_{j=0}^{k}\theta_{j}\,Q_{j}(\mathbf{x})},$
@@
3.2.5 Fourier Kolmogorov-Arnold Network (Fourier KAN)
Fourier KAN [28] uses a Fourier series expansion to capture both low- and high-frequency components in tabular or structured data. Given an input vector $\mathbf{x}\in\mathbb{R}^{d}$, the transformation function $\phi_{F}(\mathbf{x})$ introduces sine and cosine terms up to a grid size $g$, which allows the network to approximate highly complex or oscillatory functions. Formally,
$\phi_{F}(\mathbf{x})=\sum_{i=1}^{d}\sum_{k=1}^{g}\bigl{(}a_{ik}\cos(k\,x_{i})+b_{ik}\sin(k\,x_{i})\bigr{)},$
@@ 3.2.5 Fourier Kolmogorov-Arnold Network (Fourier KAN)
$y=\sum_{i=1}^{d}\sum_{k=1}^{g}\Bigl{(}W_{ik}^{(c)}\cos(k\,x_{i})+W_{ik}^{(s)}\sin(k\,x_{i})\Bigr{)}+b,$ (2)
where $W_{ik}^{(c)}$ and $W_{ik}^{(s)}$ are learnable weights for the cosine and sine terms, respectively, and $b$ is a bias. By using the orthogonality of trigonometric functions, Fourier KAN often achieves faster convergence than traditional MLPs and polynomial-based KANs while also reducing overfitting.
3.2.6 Fractional Kolmogorov-Arnold Network (fKAN)
The Fractional Kolmogorov-Arnold Network (fKAN) [29] incorporates fractional-order Jacobi functions into the Kolmogorov-Arnold framework to enhance expressiveness and adaptability. Each layer of fKAN uses a Fractional Jacobi Neural Block (fJNB), which introduces a trainable fractional parameter $\nu$ to adjust the polynomial basis dynamically. For an input $\mathbf{x}\in\mathbb{R}^{d}$, the fractional Jacobi polynomial $J_{n}^{(\alpha,\beta,\nu)}(x)$ is given by:
$J_{n}^{(\alpha,\beta)}(x^{\nu})=\frac{(\alpha+1)_{n}}{n!}\sum_{k=0}^{n}\binom{n}{k}\frac{(\beta+1)_{n-k}}{(\alpha+\beta+1)_{n-k}}\left(\frac{x^{\nu}-1}{2}\right)^{k}\left(\frac{x^{\nu}+1}{2}\right)^{n-k},$
where $(\alpha,\beta)>-1$ determine the shape of the polynomial. Within fKAN, each layer applies a linear transformation followed by a fractional Jacobi activation. This structure allows the model to capture subtle data patterns.
3.2.7 Rational Kolmogorov-Arnold Network (RKAN)